OpenStudy (anonymous):

Let f(x) = –2x + 4 and g(x) = –6x – 7. Find f(x) – g(x).

OpenStudy (anonymous):

F(x) = -2x+4; G(x) = -6x-7 F(x)-G(x)= -2x+4-(-6x) -7 = 4x-3

OpenStudy (anonymous):

That makes sense. Thank you!

OpenStudy (anonymous):

you are welcome :) ..... just combine like terms